924 Massachusetts Avenue, �2 Arlington, MA 02174

June 22, 1998

Brigham and Women’s Hospital Human Resources Attn: Mary Burns

Dear Ms. Burns,

I am writing to apply for job number 98-05-07139, senior technical research assisstant to replace Anne Brooks in Dr. Glowacki’s laboratory in Orthopedic Research. I have discussed the position with Dr. Glowacki and we have agreed that it would be possible to make the position a 4 day per week (32 hours) position rather than the posted 5 days. I have enclosed a copy of my resume for your perusal. If you have any questions please contact me at (781) 648-0367 or via email at cjsmith@mit.edu. Thank you.

Sincerely,

June 22, 1998

Tufts University Human Resources Of�ce 200 Harrison Ave. Boston, MA 02111

Dear Sir or Madam,

I am interested in any openings within your university for the position of senior research technologist or research associate. I have eight years of experience in the �elds of cellular and molecular biology, protein chemistry, and enzyme kinetics. I have recently set up and run an enzyme kinetics laboratory.

I am enclosing a copy of my resume for your perusal, and would be interested in discussing any relevant job openings. Please contact me at (781) 648-0367 or via email at cjsmith@mit.edu. Thank you for your time; I look forward to hearing from you.

Sincerely,

June 22, 1998

Director of Human Resourcs Repligen Corporation 117 Fourth Avenue Needham, MA 02194

Dear Sir or Madam,

I am interested in any openings within your university for the position of senior research technologist or research associate. I have eight years of experience in the �elds of cellular and molecular biology, protein chemistry, and enzyme kinetics. I have recently set up and run an enzyme kinetics laboratory.

I am enclosing a copy of my resume for your perusal, and would be interested in discussing any relevant job openings. Please contact me at (781) 648-0367 or via email at cjsmith@mit.edu. Thank you for your time; I look forward to hearing from you.

Sincerely,